Please help me solve for w:
7w/4 +3 = 9w/10 +6

Answer by stanbon(75887)

7w/4 +3 = 9w/10 +6
--------
Multiply thru by 20 to get::
--------
35w + 60 = 18w + 120
--------
17w = 60
----
w = 60/17
==============
Cheers,
Stan H.
==========